Putting two equal and equal charges at a distance of 3 meters, the force of repulsion of a 1.6 N between them, then what will be the value of each charge
1
Expert's answer
2020-08-27T10:25:35-0400

According to Coulomb's law, we have


"F=\frac{1}{4\pi\epsilon_0}\frac{q\cdot q}{r^2}=\frac{1}{4\pi\epsilon_0}\frac{q^2}{r^2},\\\space\\
q=2r\sqrt{\pi\epsilon_0F},\\
q=2\cdot3\sqrt{3.14\cdot8.85\cdot10^{-12}\cdot1.6}=40\cdot10^{-6}\text{ C}."

The charges are 40 μC.